# Chip-Scale Point-Source Sagnac Interferometer by Phase-Space Squeezing

### v3

@@ Results
Figure 2: Contrast and dynamic range. (A) Contrast vs. angular velocity ($\Omega$). Using 87Rb atoms, the simulation parameters are initial cloud size $\sigma_{x0}=100\,\rm{\mu m}$, temperature $T=5\,\rm{\mu K}$, and time between pulses $T_{R}=5\,\rm ms$. For the repulsive potential, we take a beam power of $P=1\,$W with a cross-section of $400\crossproduct 400\,\mu$m2 (according to the beam profile of Eq. 13), blue-detuned by $\Delta=2\pi\cdot 10\,$GHz, and focused such that in the direction of acceleration it gives rise to an inverted harmonic potential and in the transverse direction a constant potential. The harmonic profile was optimized to achieve a large squeezing factor before the atoms exceed the region where the repulsive potential is harmonic. Following Eqs. 10 and 13, and choosing $t_{\text{acc}}=80.6\,\mu$s, we find a squeezing parameter of 22.81. The purple line represents the analytical solution of PSI contrast, while the dashed orange line corresponds to SPSI, according to Eq. 5. The simulation points are the result of a numerical simulation. While the contrast reduction due to the ratio between the fringe periodicity and the initial cloud size determines the upper limit of the detection range, there exists a lower limit when the fringe period becomes larger than the final cloud size. These limits are roughly given by $\Omega_{\rm min}$ and $\Omega_{\rm max}$ (Eq. 8), which are presented in the graph as vertical dashed lines. It is evident that using the SPSI greatly improves the contrast, thereby increasing the detection dynamic range. At high angular velocities, the contrast obtained from simulation decreases faster than the analytical solution due to the short spacing between fringes relative to the detection pixel size considered only in the numerical calculation. (B) The analytical curve of the detection range as a function of $\eta$ according to Eq. 8. The dashed line presents the SPSI simulated in (A) with $\eta=22.81$.
@@ Discussion
Figure 5: Simulation fringe pattern fit. This figure compares the simulation results of both methods, the top for PSI and the bottom for SPSI. The light blue shaded area depicts the total number of atoms in each bin, denoted as $N_{\text{total}}$, while the darker shade indicates only those atoms in the internal state $|1\rangle$, denoted as $N_{|1\rangle}$. The green points represent the normalized probability for the $|1\rangle$ state, calculated as $P_{1}=N_{|1\rangle}(x)/N_{\text{total}}(x)$, while the red line denotes the sinusoidal fit to these green points. The angular velocity is determined from the fit’s fringe spatial frequency using Eq. 4. In this example the initial 1D Gaussian cloud comprising $N=10^{6}$ atoms, with a width of $\sigma_{x0}=100\,\rm{\mu m}$ and a temperature of $T=5\,\rm{\mu K}$, as in Figs. 2 and 3. For the PSI method’s simulation, the atoms undergo free expansion during the interferometer sequence, with a time interval between pulses of $T_{R}=25\,\rm ms$ used in this figure, while the simulated angular velocity is $\Omega=40\,\rm{mrad/s}$. In the case of SPSI, the cloud is additionally subjected to a repulsive potential having a power of $P=1\,\rm{W}$, with a beam cross-section of $400\crossproduct 400\,\mu$m2, as in Figs. 2 and 3, activated for $t_{\text{acc}}=20\,\rm{\mu s}$. This results in a potential frequency of $\omega=2\pi\times 1.214\,\rm{kHz}$ (Eq. 13), a squeezing parameter $\eta=5.4$ (Eq. 10), and an effective time $t_{\text{eff}}=0.84\,\rm{ms}$ (Eq. 11). Consequently, the cloud size increases notably, and we chose the potential parameters here such that it would be easy to compare to the PSI simulation. Note that the contrast and number of fringes in the PSI method is smaller than that of the SPSI (see Figs. 1-2), leading to a fit deviation and uncertainty one order of magnitude higher for the PSI case. It is important to note that while the histograms appear different in scale, the total atom count remains constant for both methods. The discrepancy in appearance is due to different bin sizes used in the PSI and SPSI plots to accommodate the varying cloud expansions, with SPSI utilizing larger bins to cover its greater spatial extent.
@@ Acknowledgments
This work was partly supported by the Israel Science Foundation Grants No. 858/18, 1314/19 and 3515/20.
@@
An unavoidable effect of the repulsive potential is the heating of the atoms induced by scattered photons (?). In the longitudinal direction of propagation of the light beam, the atoms experience an average heating of $34E_{\text{rec}}/$ 3 per scattering event, while in the two transverse directions, the heating amounts to $E_{\text{rec}}/3$ per scattering process, where $E_{\text{rec}}$ denotes the recoil energy, determining the recoil velocity $v_{\text{rec}}=\sqrt{2E_{\text{rec}}/m}$. Thus, the velocity increment in each axis for every scattering event equals $v_{\text{rec}}/\sqrt{3}$. The heating process along the acceleration direction contributes to an enlargement of the effective initial cloud’s spatial width, denoted as $\tilde{\sigma}_{x0,\text{eff}}=\sqrt{(\sigma_{x0}/\eta)^{2}+(t_{\text{eff}}\cdot\sigma_{v0,\text{heat}})^{2}}$, where $\sigma_{v0,\text{heat}}=\sqrt{N_{\text{sc}}/3}\cdot v_{\text{rec}}$ represents the heating due to the scattered photons from the repulsive potential. Here, $N_{\text{sc}}=\Gamma_{\text{sc}}t_{\text{acc}}$ signifies the number of scattering events, with $\Gamma_{\text{sc}}=(3\pi\Gamma^{2}I(\textbf{r}))/(2\hbar k_{0}^{3}c\Delta^{2})$ denoting the scattering rate. In the SPSI method, the effective reduction of $\sigma_{x0}$ augments $\Omega_{\text{max}}$, and this heating mechanism solely diminishes its value without compromising the sensitivity within the new detection range. If $(\sigma_{x0}/\eta)^{2}\gg(t_{\text{eff}}\cdot\sigma_{v0,\text{heat}})^{2}$, the effect of the scattering on the phase-space distribution becomes negligible. For the repulsive potential parameters used in Fig. 2, the relation is $(\sigma_{x0}/\eta)^{2}/(t_{\text{eff}}\cdot\sigma_{v0,\text{heat}})^{2}\approx 1$, and therefore in this example, heating only decreases $\Omega_{\text{max}}$ by a factor of $\approx\sqrt{2}$. A more detailed derivation will be provided in the following section.
@@
which is a distribution that evolves under free propagation from an uncorrelated initial distribution at $t=0$ with a spatial uncertainty $\sigma_{x0}$ and velocity uncertainty $\sigma_{v0}$. This implies that any Gaussian distribution in phase space can be obtained by free propagation over an effective time $t_{\rm eff}$ starting from an uncorrelated distribution. This is the way that the effective squeezed distribution with a squeezing factor $\eta$ is obtained from the distribution that evolves under the repulsive inverse harmonic potential after acceleration over a time $t_{\rm acc}$. It is easy to acknowledge these properties by noting that a Gaussian distribution in phase space has an elliptical shape in the $x-v$ plane, and any ellipse can be transformed into an ellipse whose axes are aligned along the main axes by a proper rotation transformation.
@@
where $n$ is the number of nodes in the function and runs from 00 to $\infty$ and $w_{n}\propto\exp(-n\hbar\omega/k_{B}T)$ are the weights (probabilities) of the states $|n\rangle$. The wave function of each of these states is a product of a Gaussian with the same width $\sigma$ for all $n$ and a Hermite polynomial and the evolution of these wave functions in free space follows the same scaling as the state $|n=0\rangle$ analyzed above. A thermal Bolzmann distribution yields a Gaussian spatial distribution $\sum_{n}w_{n}|\Phi_{n}(x)|^{2}\propto\exp(-x^{2}/2\sigma_{T}^{2})$ with $\sigma_{T}=\sqrt{k_{B}T/m\omega^{2}}$ in a harmonic trap. It can be shown that the signal resulting from an interferometric sequence that uses a thermal cloud is then identical to the signal resulting from the coherent wave packet, with the position and momentum widths $\sigma_{x0}$ and $\sigma_{v0}$ replaced by the thermal phase-space widths. These results agree completely with the results that follow from the semi-classical derivation that was used in the main text.

### v4

@@ Results
[…]
@@ Discussion
[…]
@@ Acknowledgments
This work was partly supported by the Israel Science Foundation Grants No. 858/18, 1314/19, 3515/20 and 3470/21.
@@
An unavoidable effect of the repulsive potential is the heating of the atoms induced by scattered photons (?). In the longitudinal direction of propagation of the light beam, the atoms experience an average heating of $4E_{\text{rec}}/3$ per scattering event, while in the two transverse directions, the heating amounts to $E_{\text{rec}}/3$ per scattering process, where $E_{\text{rec}}$ denotes the recoil energy, determining the recoil velocity $v_{\text{rec}}=\sqrt{2E_{\text{rec}}/m}$. Thus, the velocity increment in each axis for every scattering event equals $v_{\text{rec}}/\sqrt{3}$. The heating process along the acceleration direction contributes to an enlargement of the effective initial cloud’s spatial width, denoted as $\tilde{\sigma}_{x0,\text{eff}}=\sqrt{(\sigma_{x0}/\eta)^{2}+(t_{\text{eff}}\cdot\sigma_{v0,\text{heat}})^{2}}$, where $\sigma_{v0,\text{heat}}=\sqrt{N_{\text{sc}}/3}\cdot v_{\text{rec}}$ represents the heating due to the scattered photons from the repulsive potential. Here, $N_{\text{sc}}=\Gamma_{\text{sc}}t_{\text{acc}}$ signifies the number of scattering events, with $\Gamma_{\text{sc}}=(3\pi\Gamma^{2}I(\textbf{r}))/(2\hbar k_{0}^{3}c\Delta^{2})$ denoting the scattering rate. In the SPSI method, the effective reduction of $\sigma_{x0}$ augments $\Omega_{\text{max}}$, and this heating mechanism solely diminishes its value without compromising the sensitivity within the new detection range. If $(\sigma_{x0}/\eta)^{2}\gg(t_{\text{eff}}\cdot\sigma_{v0,\text{heat}})^{2}$, the effect of the scattering on the phase-space distribution becomes negligible. For the repulsive potential parameters used in Fig. 2, the relation is $(\sigma_{x0}/\eta)^{2}/(t_{\text{eff}}\cdot\sigma_{v0,\text{heat}})^{2}\approx 1$, and therefore in this example, heating only decreases $\Omega_{\text{max}}$ by a factor of $\approx\sqrt{2}$. A more detailed derivation will be provided in the following section.
@@
which is a distribution that evolves under free propagation from an uncorrelated initial distribution at $t=0$ with a spatial uncertainty $\sigma_{x0}$ and velocity uncertainty $\sigma_{v0}$. This implies that any Gaussian distribution in phase space can be obtained by free propagation over an effective time $t_{\rm eff}$ starting from an uncorrelated distribution. This is the way that the effective squeezed distribution with a squeezing factor $\eta$ is obtained from the distribution that evolves under the repulsive inverse harmonic potential after acceleration over a time $t_{\rm acc}$. It is easy to acknowledge these properties by noting that a Gaussian distribution in phase space has an elliptical shape in the xvx-v plane, and any ellipse can be transformed into an ellipse whose axes are aligned along the main axes by a proper rotation transformation.
@@
where $n$ is the number of nodes in the function and runs from 0 to $\infty$ and $w_{n}\propto\exp(-n\hbar\omega/k_{B}T)$ are the weights (probabilities) of the states $|n\rangle$. The wave function of each of these states is a product of a Gaussian with the same width $\sigma$ for all $n$ and a Hermite polynomial and the evolution of these wave functions in free space follows the same scaling as the state $|n=0\rangle$ analyzed above. A thermal Bolzmann distribution yields a Gaussian spatial distribution $\sum_{n}w_{n}|\Phi_{n}(x)|^{2}\propto\exp(-x^{2}/2\sigma_{T}^{2})$ with $\sigma_{T}=\sqrt{k_{B}T/m\omega^{2}}$ in a harmonic trap. It can be shown that the signal resulting from an interferometric sequence that uses a thermal cloud is then identical to the signal resulting from the coherent wave packet, with the position and momentum widths $\sigma_{x0}$ and $\sigma_{v0}$ replaced by the thermal phase-space widths. These results agree completely with the results that follow from the semi-classical derivation that was used in the main text.